Eric Poppovich invests money in two stocks for the upcoming year. His investment is shown below:

STOCK: # of shares Price Per Share Expected Return Standard Deviation
LBJ Corporation (A) 48.00 $36.31 5.00% 20.00%
Duncan Inc. (B) 29.00 $50.49 9.00% 22.00%


The correlation between LBJ and Duncan is 0.50.

What is the expected return for the portfolio in the coming year?
Answer #1

wa = (48 * 36.31) / (48*36.31 + 29 * 50.49) = 0.5434459276

wb = (29 * 50.49) / (48*36.31 + 29 * 50.49) = 0.4565540724

Expected return = 0.0682621629 = 6.82621629%